OpenStudy (anonymous):

if A+B=180(degrees) , then which of the following is true? ( 1 ) sinA=sinB ( 2 ) cosA=cosB ( 3 ) tanA=tanB a. (1) only b. (2)only c. (3)only d. all of them

hartnn (hartnn):

sin(180-x)=sin x cos(180-x)=-cos x tan(180-x)=-tan x does this help ?

OpenStudy (anonymous):

how actually will i do that?

hartnn (hartnn):

A+B=180 A=180-B

OpenStudy (anonymous):

owh yeah.. i also did that but when i tried to use 90 degrees, it went undefined when finding its tan.. so (1) and (2) are the only correct answers but it don't have that kind of answer.. will i substitute other value for A and B?? but i taught i will find myself doing it so long just to prove the equality.. :((

hartnn (hartnn):

sin(180-B)=sinB<---true cos(180-B)=cos B<---false tan(180-B)=tanB<---false

OpenStudy (anonymous):

how is that?? i dont get it :((

OpenStudy (anonymous):

owh i get it now :)) thanks!

hartnn (hartnn):

these are identities: sin(180-x)=sin x cos(180-x)=-cos x tan(180-x)=-tan x now put x=B there. also A=180-B

hartnn (hartnn):

oh, u got. great :)

OpenStudy (anonymous):

ok ok.. i forgot about the odd-even identities :)) hehe.. thanks by the way... more powers to come!! :))

hartnn (hartnn):

welcome :)